OpenStudy (anonymous):

Your fixed expenses are 1,500.45/month. Your emergency fund has 4 month's worth of coverage. You invest half in a savings account with an interest rate of 3.15% APR and the other half in a 45-day CD with an interest rate of 4.65% APR. How much is your total interest in 45 days? The answer I got was $28.85 Is that right?

OpenStudy (anonymous):

@jim_thompson5910 Feel like checking this? You're my go to man now! (:

jimthompson5910 (jim_thompson5910):

It sounds like you are paying $1,500.45 a month, for 4 months, into some fund Afterwards, the interest is accrued for 45 days -------------------- So you have paid 1500.45*4 = 6001.80 dollars total This is the principle which is to be split in half among the two investment options -------------------- You invest 6001.80/2 = 3000.90 dollars in the savings account So you gain 3000.90*(0.0315)*(45/365) = 11.6541801369863 = 11.65 dollars from the savings account alone You invest 3000.90 in the CD So you gain 3000.90*(0.0465)*(45/365) = 17.2037897260273 = 17.20 dollars from the CD alone --------------------- So in total, you get back 11.65 + 17.20 = 28.85 dollars in interest So it looks like you are correct

OpenStudy (anonymous):

Yay! Thank you. (:

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (agent0smith):

Looks good to me